# Technical insights and legal considerations for advancing federated learning in bioinformatics

### v1

@@ Abstract
Federated learning leverages data across institutions to improve clinical discovery while complying with data-sharing restrictions and protecting patient privacy. As the evolution of biobanks in genetics and systems biology has proved, accessing more extensive and varied data pools leads to a faster and more robust exploration and translation of results. More widespread use of federated learning may have the same impact in bioinformatics, allowing access to many combinations of genotypic, phenotypic and environmental information that are undercovered or not included in existing biobanks. This paper reviews the methodological, infrastructural and legal issues that academic and clinical institutions must address before implementing it. Finally, we provide recommendations for the reliable use of federated learning and its effective translation into clinical practice.
keywords:
@@
Federated Machine Learning, Exposome, Secure distributed analysis, Data privacy, Collaborative genomics1 Introduction
@@
Sharing personal information has been increasingly regulated in both the EU (with the GDPR and the AI act; European Union, 2016, 2024) and the US (with the National AI Initiative Act; U.S. Congress, 2020) to mitigate the personal and societal risks associated with their use, particularly in connection with machine learning and AI models (Cath et al., 2018). These regulations make multi-centre studies and similar endeavours more challenging, impacting biomedical and clinical research.
@@
Federated learning (FL; McMahan et al., 2017; Ludwig and Baracaldo, 2022) is a technical solution intended to reduce the impact of these restrictions. FL allows multiple parties to train a global machine learning model collaboratively from the respective data without sharing the data themselves and without any meaningful model performance degradation. Instead, parties only share model updates, making it impractical to reconstruct personal information when the appropriate secure computational measures are implemented (Wainakh et al., 2022).
@@
This approach strengthens security by keeping sensitive information local, improves privacy by minimising data exposure even between the parties involved, and limits risk of data misuse by allowing each party to retain complete control over its data (Truong et al., 2021). If enough parties are involved, FL may potentially access larger and more varied data pools than centralised biobanks can provide, particularly if there are legal (or other) barriers to data centralisation, resulting in more accurate and robust models than those produced by any individual party.
@@
FL has proven to be a valuable tool for biomedical research expected to gain further traction in the coming years. Its use has improved breast density classification models (accuracy up by 6%, generalisability up by 46%; Roth et al., 2020), COVID-19 outcome prediction at both 24h and 72h (up 16% and 38%; Dayan et al., 2021) and rare tumour segmentation (up by 23-33% and 15%; Pati et al., 2022) compared to single-party analyses. A consortium of ten pharmaceutical companies found that FL improved structure-activity relationship (QSAR) models for drug discovery (both up 12% Heyndrickx et al., 2023). Early-stage applications building predictive models from electronic health records (Brisimi et al., 2018) have also confirmed no practical performance degradation compared to pooling data from all parties.
@@
To achieve such results, a real-world implementation of FL must overcome several methodological, infrastructural and legal issues. However, biomedical FL literature reviews (Xu et al., 2020; Chowdhury et al., 2022, among others) are predominantly high-level and considered simulated rather than real-world implementations. Here, we will cover federated methods designed explicitly for bioinformatics and discuss the infrastructure they need and how they satisfy the legal requirements.
@@
To this end, we have structured the remainder of the paper as follows: We first review the fundamental concepts and design decisions of FL in Section 2, including different topologies (Section 2.1), hardware and software (Section 2.2), data layouts in different parties (Sections 2.3 and 2.4), security (Section 2.5) and privacy concerns (Section 2.6). In Section 3, we contrast and compare bioinformatics FL methods for proteomics and differential expression (Section 3.1), genome-wide association studies (GWAS; Section 3.2), single-cell RNA sequencing (Section 3.3), multiomics (Section 3.4) and medical imaging (Section 3.5) applications. We conclude the section with notable examples of ready-to-use software tools (Section 3.6). Section 4 provides examples of federated operations common in bioinformatics. Finally, we discuss the relevant legal implications of using FL (Section 5) before discussing our perspective on its reliable use and translation to clinical practice (Section 6).
@@ 2 Federated learning
FL is a collaborative approach to machine learning model training, where multiple institutions form a consortium to jointly train a shared model by exchanging model updates rather than raw patient data. Typically, FL involves data holders (called “clients”) sharing their local contributions with a server (McMahan et al., 2017) as outlined in Figure 1. The server then creates and shares back a global model, inviting the data holders to update and resubmit their contributions. This process is iterative and involves several rounds of model update exchanges. Unlike traditional centralised computing, FL does not store patient data in a central location. Instead, patient data remain under the control of the respective data owners at their sites, enhancing privacy.
@@
FL has similarities with distributed computing, meta-analysis, and trusted research environments (TREs) but also has key differences, which we highlight below.
@@
Distributed computing (DC) (Zomaya, 2006) divides a computational task among multiple machines to enhance processing speed and efficiency. Usually, DC starts from a centrally managed data set spread across different machines, which is assumed to contain independent and identically distributed observations. Each machine is tasked to process a comparable quantity of data. In contrast, clients independently join FL with their locally-held data, which may vary significantly in quantity and distribution. While sharing some techniques with FL, distributed computing aims for computational efficiency and lacks its privacy focus.
@@
On the other hand, meta-analyses (Toro-Domínguez et al., 2021) aggregate results across previously completed studies using statistical methods to account for their variations, thus allowing researchers to synthesise findings without accessing raw data and preserve the privacy of individual data sets. Here, FL collaboratively trains a joint model using distributed data to iteratively update it while meta-analysis constructs it in a single step from the pre-existing results. Multiple studies on sequencing data have demonstrated that FL produces results closer to centralised analysis than from meta-analysis (Mendelsohn et al., 2023; Zolotareva et al., 2021).
@@
TREs (Kavianpour et al., 2022) provide access to data within a controlled, secure computing environment for conducting analyses, almost always disallowing data sharing. Some TREs have a centralised data location and governance; an example is the Research Analysis Platform (RAP), the TRE for the UK Biobank (UKB; Sudlow et al., 2015). Others, such as FEGA (Federated European Genome-phenome Archive, 2024), are decentralised. Each institution maintains its data locally; only the relevant data are securely transferred to the computing environment when the analysis is authorised. Unlike FL, the learning process is not distributed across the data holders. Thus, the tradeoff between TREs and FL is between a centralised, trusted entity with large computational facilities that can place substantial restrictions on the analysis, and a consortium that requires all parties to apply governance guidelines and provide compute, but can scale both data access and privacy guarantees.
@@ 2.1 Topologies
The topology of the FL consortium is determined by the number of participating parties and their defined interactions. Some examples are illustrated in Figure 2. The most common is the centralised topology, where multiple data-holding parties (the clients) collaboratively train a shared machine learning model through a central server (the aggregator) that iteratively collects model updates from each client, updates the global model, and redistributes it back to the clients. Typically, clients do not communicate directly; they only communicate with the central server. In contrast, a decentralised topology (Beltrán et al., 2023) lacks a dedicated aggregation server. All consortium parties can potentially be model trainers and aggregators, interacting via peer-to-peer communication. Hybrid configurations include, for instance, using two servers: one server handles aggregation of noisy local models, while the other performs auxiliary tasks, such as noise aggregation (Nasirigerdeh et al., 2021). Clients can communicate with the servers, and servers can communicate between themselves, but clients cannot communicate with each other.
@@ 2.2 Hardware and software
Hardware, software and models should be chosen with knowledge of the data and inputs from domain and machine learning specialists to design an effective machine learning pipeline (Scutari and Malvestio, 2023).
In terms of infrastructure, FL requires computational resources for each client and server. The optimal hardware configuration depends on the models to be trained; at a minimum, each client must be able to produce model updates from local data, and each server must be able to aggregate those updates and manage the consortium. Connection bandwidth is not necessarily critical: to date, client-server communications contain only a few megabytes of data, reaching 150MB only for large computer vision models, and can be made more compact through compression and model quantisation (Camajori Tedeschini et al., 2022). On the other hand, latency may be a bottleneck if it limits the hardware utilisation.
As for software, several dedicated FL frameworks, many of which are comparatively analysed in (Riedel et al., 2024), provide structured tools and environments for developing, deploying, and managing federated machine learning models. While some frameworks, such as Tensorflow Federated (TFF; Google, 2024), specialise in particular models, others support a broader range of approaches. Notable open-source examples include PySyft (Ziller et al., 2021) and Flower (Beutel et al., 2020). Both are supported by active communities and integrate with PyTorch to train complex models. PySyft is a multi-language library focusing on advanced privacy-preserving techniques, including differential privacy and homomorphic encryption. Flower is an FL framework: its modular design and ease of customisation make it particularly useful for large-scale and multi-omics studies involving heterogeneous devices and clients. We will provide examples using these frameworks in Section 3 before discussing frameworks explicitly designed for bioinformatics in Section 3.6.
Other frameworks target healthcare and biomedical applications but not bioinformatics specifically. For instance, OpenFL (Foley et al., 2022) is designed to facilitate FL on sensitive EHRs and medical imaging data; it supports different data partitioning schemes (Section 2.4) but struggles with heterogeneous cross-device FL (Section 2.3). NVIDIA Clara, which was used in Dayan et al. (2021), has similar limitations.
@@ 2.3 Usage scenarios: cross-device and cross-silo
FL applications take different forms in different domains. Many small, low-powered clients, such as wearable medical devices from the Internet of Things, may produce the data needed to train the federated machine learning model. Such cross-device communications are often unreliable: passing lightweight model updates instead of raw data largely addresses connectivity issues and privacy risks.
FL may also involve a small number of parties, each possessing large amounts of sensitive data (Huang et al., 2022), stored within their “data silos”. In this cross-silo scenario, common in healthcare and bioinformatics, the priority is to minimise the privacy risks associated with data sharing and comply with regulations. Minimising large data transfers is also computationally advantageous when modelling large volumes of information, such as whole-genome sequences.
@@ 2.4 Data partitioning and heterogeneity
Data may be partitioned along two axes: each party may record the same features for different samples or features describing the same samples (Figure 3). In the first scenario, known as horizontal FL,111This naming convention assumes that samples (features) are the rows (columns) of a tabular data set. different parties may each possess genomic sequencing data from different individuals. In contrast, in vertical FL, one party may hold data from one omic type (say, genomic data), while another may have data from a different phenotype or omic type (say, proteomic data) for the same individuals. Horizontal FL is by far the most prevalent approach in bioinformatics.
Significant variations in sample size and feature distributions between data holders often exist. This heterogeneity allows FL to better capture the variability of the underlying population, resulting in transferrable models that generalise well (Sheller et al., 2020). Clearly, if data holders collect observations from distinct populations, any federated model trained from them must be correctly specified to capture population structure and avoid bias in inference and prediction. If the populations are known, we can train targeted population-specific models alongside the global one (Tan et al., 2022). Otherwise, we can use clustering to identify them from the available data (Sattler et al., 2021). Accounting for variations in measurements, definitions and distributions to harmonise data across parties is also fundamental but is much more challenging because access to data is restricted, even more so than in meta-analysis (Camajori Tedeschini et al., 2022).
@@ 2.5 Security and privacy
In terms of privacy, deep learning models are the most problematic in machine learning because of their ability to memorise training data. They leak individual observations during training (through model updates; Geiping et al., 2020), after training (through their parameters; Haim et al., 2022) and during inference (membership attacks; Shokri et al., 2017; Hu et al., 2022). More broadly, individual reidentification is an issue for genetic data (Homer et al., 2008) and all the models learned from them. For instance, (Cai et al., 2015) has demonstrated that it is possible to identify an individual from the linear model learned in an association study from just 25 genes. However, such works make unrealistic assumptions on the level of access to the models and the data (Kolobkov et al., 2024): even basic infrastructure security measures and the distributed nature of the data will make such identification difficult under the best circumstances. The privacy-enhancing techniques discussed in Section 2.6 can make such efforts completely impractical.
As for security, we must consider different threat models, understanding what information requires protection, their vulnerabilities, and how to mitigate or respond to threats. Internal and external threats to the consortium should be treated equally with security in depth design and implementation decisions that consider parties untrusted. Security threats, such as membership attacks and model inversion attacks (Fredrikson et al., 2015), can originate equally from parties and external adversaries that seek to abuse the model inference capabilities to extract information about the data. On the other hand, adversarial attacks are more likely to originate from consortium parties that seek to introduce carefully crafted data or model updates into the training process to produce a global model with undesirable behaviour. Some examples are data poisoning (Sun et al., 2022b), manipulation (Jagielski et al., 2018) and Byzantine attacks (Li et al., 2023a).
Encrypting communication channels, implementing strict authentication (to verify each party’s identity) and authorisation (to control which information and resources each party has access to or shares) schemes, and keeping comprehensive access logs for audit can secure any machine learning pipeline, including federated ones. Similarly, using an experiment tracking platform makes it possible to track data provenance, audit both the data and the training process and ensure the reproducibility of results (Scutari and Malvestio, 2023). These measures must be complemented by federated models resistant to these threats at training and inference time, as thoroughly discussed in Yin et al. (2021).

### v4

# Technical and legal aspects of federated learning in bioinformatics: applications, challenges and opportunities

@@ Abstract
Federated learning leverages data across institutions to improve clinical discovery while complying with data-sharing restrictions and protecting patient privacy. This paper provides a gentle introduction to this approach in bioinformatics, and is the first to review key applications in proteomics, genome-wide association studies (GWAS), single-cell and multi-omics studies in their legal as well as methodological and infrastructural challenges. As the evolution of biobanks in genetics and systems biology has proved, accessing more extensive and varied data pools leads to a faster and more robust exploration and translation of results. More widespread use of federated learning may have a similar impact in bioinformatics, allowing academic and clinical institutions to access many combinations of genotypic, phenotypic and environmental information that are undercovered or not included in existing biobanks.
[…]
@@
Federated machine learning, Exposome, Secure distributed analysis, Data privacy, Collaborative genomics
2 Introduction
@@
Sharing personal information has been increasingly regulated in both the EU (with the GDPR and the AI act; 1, 2) and the US (with HIPAA and the National AI Initiative Act; 3, 4) to mitigate the personal and societal risks associated with their use, particularly in connection with machine learning and AI models (5). These regulations make multi-centre studies and similar endeavours more challenging, impacting biomedical and clinical research.
@@
Federated learning (FL; 6, 7) is a technical solution intended to reduce the impact of these restrictions. FL allows multiple parties to collaboratively train a global machine learning model using their respective data without sharing it themselves, and without any meaningful model performance degradation. Instead, parties only share model updates, making it impractical to reconstruct personal information when the appropriate secure computational measures are implemented (8).
@@
This approach strengthens security by keeping sensitive information local, improves privacy by minimising data exposure even between the parties involved, and limits risk of data misuse by allowing each party to retain complete control over its data (9). If enough parties are involved, FL may access larger and more varied data pools than centralised biobanks can provide. This is particularly true if there are legal (or other) barriers to data centralisation, resulting in more accurate and robust models than those produced by any individual party.
@@
FL has proven to be a valuable tool for biomedical research and is expected to gain further traction in the years to come. Its use has improved breast density classification models (accuracy up by 6%, generalisability up by 46%; 10), COVID-19 outcome prediction at both 24h and 72h (up 16% and 38%; 11) and rare tumour segmentation (up by 23-33% and 15%; 12) compared to single-party analyses. A consortium of ten pharmaceutical companies found that FL improved structure-activity relationship (QSAR) models for drug discovery (both up 12% 13). Early-stage applications building predictive models from electronic health records (14) have also confirmed no practical performance degradation compared to pooling data from all parties.
@@
To achieve such results, a real-world implementation of FL must overcome several methodological, infrastructural and legal issues. However, biomedical FL literature reviews (15, 16, among others) are predominantly high-level and considered simulated rather than real-world implementations. Here, we will cover federated methods designed explicitly for bioinformatics and discuss the infrastructure they require, as well as how they meet legal requirements. In reviewing the literature, we selected papers that study practical analysis problems (as opposed to proposing methodologies in the abstract) for proteomics, genome-wide association studies (GWAS), and single-cell and multi-omics data. We also considered papers that discuss their feasibility, trade-offs, and performance compared to centralised analyses, and were published after 2016. We used Google Scholar to find and retrieve them.
@@
To this end, we have structured the remainder of the paper as follows: We first review the fundamental concepts and design decisions of FL in Section 3, including different topologies (Section 3.1), hardware and software (Section 3.2), data layouts in different parties (Sections 3.3 and 3.4), security (Section 3.5) and privacy concerns (Section 3.6). In Section 4, we contrast and compare bioinformatics FL methods for proteomics and differential expression (Section 4.1), genome-wide association studies (GWAS; Section 4.2), single-cell RNA sequencing (Section 4.3), multiomics (Section 4.4) and medical imaging (Section 4.5) applications. We conclude the section with notable examples of ready-to-use software tools (Section 4.6). Section 5 provides examples of federated operations common in bioinformatics. Finally, we discuss the legal implications of using FL (Section 6) before summarising our perspective in Section 8.
@@ 3 Federated learning
FL is a collaborative approach to machine learning model training, where multiple institutions form a consortium to jointly train a shared model by exchanging model updates rather than individual patient data. Typically, FL involves data holders (called ”clients”) sharing their local contributions with a server (6) as outlined in Figure 1. The server then creates and shares back a global model, inviting the data holders to update and resubmit their contributions. This process is iterative and involves several rounds of model update exchanges. Unlike traditional centralised computing, FL does not store patient data in a central location. Instead, patient data remain under the control of the respective data owners at their sites, enhancing privacy.
@@
FL has similarities with distributed computing, meta-analysis, and trusted research environments (TREs), but also has key differences, which we highlight below. Table 1 provides a comparative overview of these approaches.
@@
Distributed computing (DC) (17) divides a computational task among multiple machines to enhance processing speed and efficiency. Typically, DC starts from a centrally managed data set spread across multiple machines, which is assumed to contain independent and identically distributed observations. Each machine is tasked to process a comparable quantity of data. In contrast, clients independently join FL with their locally held data, which may vary significantly in quantity and distribution. While sharing some techniques with FL, distributed computing aims for computational efficiency and lacks its privacy focus.
@@
On the other hand, meta-analyses (18) aggregate results across previously completed studies using statistical methods to account for their variations, thus allowing researchers to synthesise findings without accessing personal data and preserve the privacy of individual data sets. Here, FL collaboratively trains a joint model using distributed data to iteratively update it while meta-analysis constructs it in a single step from the pre-existing results. Multiple studies on sequencing data have demonstrated that FL produces results closer to centralised analysis than from meta-analysis (19, 20).
@@
TREs (21) provide access to data within a controlled, secure computing environment for conducting analyses, almost always disallowing data sharing. Some TREs have a centralised data location and governance; an example is the Research Analysis Platform (RAP), the TRE for the UK Biobank (UKB; 22). Others, such as FEGA (23), are decentralised. Each institution maintains its data locally; only the relevant data are securely transferred to the computing environment when the analysis is authorised. Unlike FL, the learning process is not distributed across the data holders. Thus, the trade-off between TREs and FL is between a centralised, trusted entity with extensive computational facilities that can place substantial restrictions on the analysis, and a consortium that requires all parties to apply governance guidelines and provide compute, but can scale both data access and privacy guarantees.
| Aspect | Centralised Learning | Federated Learning | Distributed Computing | Meta-analysis | Trusted Research Environments (TRE) |
| Primary goal | Aggregate all individual data into one place and train or analyse centrally. | Collaborative model training across parties without sharing individual data. | Increase speed and scalability; job parallelisation. | Combine evidence from completed studies. | Provide secure, auditable access to sensitive data for research. |
| Where individual data live | Single central repository. | Stay local at each device/institution. | Centrally stored, sharded across nodes. | Remain with original studies; not pooled. | In a secure environment or under local (federated) control (only relevant data are transferred). |
| How learning happens | Training/analysis is run on pooled data in one environment. | Participants compute local model updates and send them for secure aggregation in iterative rounds. | Tasks are partitioned and executed in parallel; results are combined centrally. | Study-level results are aggregated. | Researchers run code/queries inside the TRE; outputs are checked before release. |
| Participation | All data contributors must share data with the central site beforehand. | Multiple data holders, dynamic participation possible (devices can join/leave). | Centrally managed workers/nodes with data partitions. | Fixed set of completed/published studies. | Approved users/projects with strict governance and access control. |
| Data assumptions | No inherent assumption; depends on chosen analysis method. | Must handle non-IID data and uneven sample sizes. | Often assumes roughly IID, evenly partitioned data. | Models between-study heterogeneity (fixed/random effects). | No inherent assumption; depends on chosen analysis method. |
| What moves across parties | Individual data sent to the central site. | Model updates (gradients/weights), possibly in shares (SMPC), encrypted or differentially private. | Data blocks and intermediate results. | Study-level summary statistics. | Code/queries go in; vetted results come out. |
| Privacy posture | Highest data exposure (requires trust in central data custodian). | Designed to avoid individual data sharing; can support privacy- enhancing techniques. | Not privacy-focused (single trust domain). | Only summary results shared. | Via technical or organisational controls. |
| Output artefact | Single trained model or analysis result from pooled data. | Global or personalised model held by each participant. | Finished job outputs. | Summary results with uncertainty estimates. | Analysis outputs are released after disclosure control. |
| Typical examples | Central data warehouse, pooled data in a multi-centre study. | Cross-hospital FL; edge device FL. | Spark, Hadoop, Ray, HPC clusters. | Cochrane-style meta-analyses. | UK Biobank RAP, Federated EGA. |
| Legal responsibilities | The central data controller has the responsibility for legal compliance and security. | Data controllers retain responsibility for legal compliance and security; data processors have contractual responsibilities linked to that. | Depends on data origin: same as centralised learning for single-centre studies, or as federated learning when data comes from multiple centres. | Data controllers retain responsibility; data processors must ensure original data use agreements permit meta-analysis. | The operator is responsible for TRE security and governance. Data controllers retain legal responsibility for sharing the data. |
| Legal basis in addition to data subjects’ consent | Only data subjects’ informed consent is needed. | Data sharing agreements for pseudo-anonymised data. | Data sharing agreement for individual data in case data from multiple centres are aggregated. | No personal data involved if the data are sufficiently aggregated (anonymised); otherwise, same as federated learning. | Access agreements between TREs and data controllers: permitted uses, audit, security protocols. Data processors’ agreements with TRE. |
@@ 3.1 Topologies
The topology of the FL consortium is determined by the number of participating parties and their defined interactions. Some examples are illustrated in Figure 2. The most common is the centralised topology, where multiple data-holding parties (the clients) collaboratively train a shared machine learning model through a central server (the aggregator) that iteratively collects model updates from each client, updates the global model, and redistributes it back to the clients. Typically, clients do not communicate directly; they only communicate with the central server. In contrast, a decentralised topology (24) lacks a dedicated aggregation server. All consortium parties can potentially serve as model trainers and aggregators, interacting through peer-to-peer communication. Hybrid configurations include, for instance, using two servers: one server handles aggregation of noisy local models, while the other performs auxiliary tasks, such as noise aggregation (25). Clients can communicate with the servers, and servers can communicate with each other, but clients cannot communicate with each other.
@@ 3.2 Hardware and software
Hardware, software and models should be chosen with knowledge of the data and inputs from domain and machine learning specialists to design an effective machine learning pipeline (26).
In terms of infrastructure, FL requires computational resources for each client and server. The optimal hardware configuration depends on the models to be trained; at a minimum, each client must be able to produce model updates from local data, and each server must be able to aggregate those updates and manage the consortium. Connection bandwidth is not necessarily critical: to date, client-server communications contain only a few megabytes of data, reaching 150MB only for large computer vision models, and can be made more compact through compression and model quantisation (27). On the other hand, latency may be a bottleneck if it limits the hardware utilisation.
As for software, several dedicated FL frameworks, many of which are comparatively analysed in (28), provide structured tools and environments for developing, deploying, and managing federated machine learning models. While some frameworks, such as Tensorflow Federated (TFF; 29), specialise in particular models, others support a broader range of approaches. Notable open-source examples include PySyft (30) and Flower (31). Both are supported by active communities and integrate with PyTorch to train complex models. PySyft is a multi-language library focusing on advanced privacy-preserving techniques, including differential privacy and homomorphic encryption. Flower is an FL framework: its modular design and ease of customisation make it particularly useful for large-scale and multi-omics studies involving heterogeneous devices and clients. We will provide examples using these frameworks in Section 4 before discussing frameworks explicitly designed for bioinformatics in Section 4.6.
Other frameworks target healthcare and biomedical applications, but not bioinformatics specifically. For instance, OpenFL (32) is designed to facilitate FL on sensitive EHRs and medical imaging data; it supports different data partitioning schemes (Section 3.4) but struggles with heterogeneous cross-device FL (Section 3.3). NVIDIA Clara, which was used in Dayan et al. (11), has similar limitations.
@@ 3.3 Usage scenarios: cross-device and cross-silo
FL applications take different forms in different domains. Many small, low-powered clients, such as wearable medical devices from the Internet of Things, may produce the data needed to train the federated machine learning model. Such cross-device communications are often unreliable: passing lightweight model updates instead of individual data largely addresses connectivity issues and privacy risks.
FL may also involve a small number of parties, each possessing large amounts of sensitive data (33), stored within their ”data silos”. This setting, often called the emphcross-silo scenario, is common in healthcare and bioinformatics. Here, the main priority is to minimise the privacy risks associated with data sharing and comply with regulations. Additionally, minimising large data transfers is also computationally advantageous when modelling large volumes of information, such as whole-genome sequences.
@@ 3.4 Data partitioning and heterogeneity
Data may be partitioned along two axes: each party may record the same features for different samples or features describing the same samples (Figure 3). In the first scenario, known as horizontal FL, different parties may each possess genomic sequencing data from different individuals. In contrast, in vertical FL, one party may hold data from one omic type (say, genomic data), while another may have data from a different phenotype or omic type (say, proteomic data) for the same individuals. Horizontal FL is by far the most prevalent approach in bioinformatics.
Significant variations in sample size and feature distributions between data holders often exist. This heterogeneity allows FL to better capture the variability of the underlying population, resulting in transferable models that generalise well (34). Clearly, if data holders collect observations from distinct populations, any federated model trained from them must be correctly specified to capture population structure and avoid bias in inference and prediction. If the populations are known, we can train targeted population-specific models alongside the global one (35). Otherwise, we can use clustering to identify them from the available data (36). Accounting for variations in measurements, definitions and distributions to harmonise data across parties is also fundamental but is much more challenging because access to data is restricted, even more so than in meta-analysis (27).
@@ 3.5 Security and privacy
In terms of privacy, deep learning models are the most problematic in machine learning because of their ability to memorise training data. They leak individual observations during training (through model updates; 37), after training (through their parameters; 38) and during inference (membership attacks; 39, 40). More broadly, individual reidentification is an issue for genetic data (41) and all the models learned from them. For instance, (42) has demonstrated that it is possible to identify an individual from the linear model learned in an association study from just 25 genes. However, such works make unrealistic assumptions on the level of access to the models and the data (8): even basic infrastructure security measures and the distributed nature of the data will make such identification difficult under the best circumstances. The privacy-enhancing techniques discussed in Section 3.6 can make such efforts wholly impractical.
As for security, we must consider different threat models, understanding what information requires protection, their vulnerabilities, and how to mitigate or respond to threats. Internal and external threats to the consortium should be treated equally with security in depth design and implementation decisions that consider parties untrusted. Security threats, such as membership attacks and model inversion attacks (43), can originate equally from parties and external adversaries that seek to abuse the model inference capabilities to extract information about the data. On the other hand, adversarial attacks are more likely to originate from consortium parties that seek to introduce carefully crafted data or model updates into the training process to produce a global model with undesirable behaviour. Some examples are data poisoning (44), manipulation (45) and Byzantine attacks (46).
Encrypting communication channels, implementing strict authentication (to verify each party’s identity) and authorisation (to control which information and resources each party has access to or shares) schemes, and keeping comprehensive access logs for audit can secure any machine learning pipeline, including federated ones. Similarly, using an experiment tracking platform makes it possible to track data provenance, audit both the data and the training process and ensure the reproducibility of results (26). These measures must be complemented by federated models resistant to these threats at training and inference time, as thoroughly discussed in Yin et al. (47).